OpenStudy (anonymous):

if the period of a cosine function is 7pi/4 what is the frequency of the function?

OpenStudy (anonymous):

flip period , you will get frequency

OpenStudy (anonymous):

no that's not one of the answers.

OpenStudy (amistre64):

the process isnt one of your options?

OpenStudy (anonymous):

the choices are 8/7 8pi/7 7/8 pi

OpenStudy (amistre64):

ok, but what defines frequency?

OpenStudy (amistre64):

The frequency is the reciprocal of the period. http://www.algebralab.org/lessons/lesson.aspx?file=trigonometry_trigperiodfreq.xml

OpenStudy (mertsj):

So perhaps you have posted the problem incorrectly or got the answer choices mixed up with another problem.

OpenStudy (amistre64):

how many given periods are in 2pi ? is what im thinking the frequency would measure

OpenStudy (amistre64):

2pi/(7pi/4) = 4(2/7) = 8/7

OpenStudy (anonymous):

oh i see, thanks